# CLQCD Collaboration

### v2

@@ IV Summary
Our central finding is the observation of a favorable $\mathcal{O}(a^{4})$ scaling for $\Delta_{\text{mix}}$ when valence overlap or clover fermions are used on gauge ensembles generated with 2+1+1 flavor HISQ sea fermions. This result is particularly significant as it was achieved using the same tadpole-improved Symanzik gauge action previously associated with a sizable $\Delta_{\text{mix}}$ in CLQCD clover fermion ensembles Zhao:2022ooq. This comparison directly verifies that the use of a dynamical fermion action with chiral symmetry is the dominant factor in suppressing mixed-action artifacts, irrespective of the gauge action. Furthermore, by comparing different action combinations at a fixed lattice spacing, we have quantified additional, secondary influences on $\Delta_{\text{mix}}$. Our results show that the larger correction from rectangle loops in the gauge action can provide an additional suppression of $\Delta_{\text{mix}}$, while the effect of dynamical charm quarks is statistically insignificant.
Our finds here offers increased flexibility in selecting valence fermion actions for a given set of gauge configurations, thereby enhancing the potential of the mixed-action approach. Further investiagation on different choice of the staggered fermion (likes the stout smearing with different steps and smearing sizes) would be helpful to further suppress $\Delta_{\text{mix}}$ and worth studies in the future.
Furthermore, a preliminary study lin2025 of physical observable using clover valence fermions on our HISQ ensembles indicates that the discretization errors for quantities such as $f_{K}$ and $m_{\Omega}$ are suppressed compared to those from unitary clover fermion calculations CLQCD:2023sdb; Hu:2024mas. This suggests that the mixed-action setup can offer superior control over systematic uncertainty from continuum extrapolation, compared to the one with unitary valence-sea fermion action. Consequently, further investigations into a wider range of observables using this setup are warranted to identify optimal, cost-efficient strategies for lattice QCD calculations that maintain rigorous control over systematic errors.
@@ Acknowledgement
We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions CLQCD:2023sdb, which are generated on HPC Cluster of ITP-CAS, IHEP-CAS and CSNS-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. The calculations were performed using the PyQUDA package Jiang:2024lto with QUDA Clark:2009wm; Babich:2011np; Clark:2016rdz through HIP programming model Bi:2020wpt. The numerical calculation were carried out on the ORISE Supercomputer, HPC Cluster of ITP-CAS and Advanced Computing East China Sub-center. This work is supported in part by National Key R&D Program of China No.2024YFE0109800, NSFC grants No. 12435002, 12525504, 12293060, 12293062, and 12447101, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. YSBR-101 and XDB34030303, and the science and education integration young faculty project of University of Chinese Academy of Sciences.
@@ IV.1 Autocorrelation Analysis
To assess the statistical independence of gauge configurations generated by the Hybrid Monte Carlo algorithm, we perform an autocorrelation analysis using the binning method following the approach in Ref. Du:2024. The analysis focuses on four key observables: the average plaquette $\langle\text{Tr }U_{p}\rangle/3$, the topological charge $Q=\int d^{4}xq(x)$ with $q(x)=\frac{g^{2}}{32\pi^{2}}\epsilon_{\mu\nu\rho\sigma},\text{Tr},[F^{\mu\nu}(x)F^{\rho\sigma}(x)]$, the pion mass $m_{\pi}$, and the $\eta_{c}$ mass $m_{\eta_{c}}$.
For a given observable $O$, the normalized variance ratio $\sigma_{n}/\sigma_{1}$ is computed as a function of bin size $n$. First, the overall mean is calculated as $\bar{O}=\frac{1}{N}\sum_{i=1}^{N}O_{i}$, where $O_{i}$ denotes an individual measurement and $N$ is the total number of configurations. For a bin size $n$, the data is partitioned into $M=\left\lfloor\frac{N}{n}\right\rfloor$ bins, and the mean for the $k$-th bin is given by $\bar{O}{n,k}=\frac{1}{n}\sum{i=(k-1)n+1}^{kn}O_{i}$. The variance of these binned means is then computed as $\sigma_{n}^{2}(O)=\frac{1}{M-1}\sum_{k=1}^{M}(\bar{O}{n,k}-\bar{O})^{2}$. To estimate the uncertainty of this variance, we calculate its standard error, $\mathrm{SE}(\sigma_{n}^{2})=\sqrt{\frac{1}{M(M-1)}\sum{k=1}^{M}\left(s_{n,k}^{2}-\sigma_{n}^{2}\right)^{2}}$, where $s_{n,k}^{2}(O)=\frac{1}{n-1}\sum_{i=(k-1)n+1}^{kn}(O_{i}-\bar{O}_{n,k})^{2}$ is the variance within the $k$-th bin.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
The left three panels of Fig. 5 illustrate the ratio $\sigma_{n}/\sigma_{1}$ as a function of bin size $n$ for all four observables, using data from the ensemble with $m_{\pi}\sim 0.3$ GeV at $a=0.108$ fm. These panels correspond to different molecular dynamics time steps: $\tau=0.5$ (upper left), $1.0$ (central left), and $2.0$ (lower left). As $\tau$ increases, the dependence of the statistical error on bin size weakens for all quantities. The three panels on the right present the cases for ensembles at finer lattice spacings ($a=0.09$, $0.07$, and $0.05$ fm from top to bottom), all with $\tau=2$. These plots suggest that the sensitivity to binning does not become significantly stronger at smaller lattice spacings.

### v3

@@ IV Summary
Our central finding is the observation of a favorable $\mathcal{O}(a^{4})$ scaling for $\Delta_{\text{mix}}$ when valence overlap or clover fermions are used on gauge ensembles generated with 2+1+1 flavor HISQ sea fermions. This result is particularly significant as it was achieved using the same tadpole-improved Symanzik gauge action previously associated with a sizable $\Delta_{\text{mix}}$ in CLQCD clover fermion ensembles Zhao et al. (2023). This comparison directly verifies that the use of a dynamical fermion action with chiral symmetry is the dominant factor in suppressing mixed-action artifacts, irrespective of the gauge action. Furthermore, by comparing different action combinations at a fixed lattice spacing, we have quantified additional, secondary influences on $\Delta_{\text{mix}}$. Our results show that the larger correction from rectangle loops in the gauge action can provide an additional suppression of $\Delta_{\text{mix}}$, while the effect of dynamical charm quarks is statistically insignificant.
A further comparison with values from the literature Cichy et al. (2013); Orginos and Walker-Loud (2008); Aubin et al. (2008); Walker-Loud and others (2009) indicates that using either twisted mass or AS fermions as the sea action does not reduce mixed-action effects compared to the HISQ fermion. For the valence fermion, the domain wall fermion yields a $\Delta_{\text{mix}}$ Berkowitz and others (2017) comparable to that of either the clover or overlap fermion. If clover valence fermions are used, HYP smearing is preferable. The Borici-Creutz fermion appears to be a promising candidate for the valence fermion action Basak et al. (2017), although higher-precision calculations at finer lattice spacings are required to confirm this.
In summary, our studies here that the HISQ sea fermion with the Iwasaki gauge action is the so far most cost-effective combination for suppressing $\Delta_{\text{mix}}$ with the flexibility in selecting valence fermion actions. If clover valence fermions are used, HYP smearing is preferable. The tadpole-improved Symanzik gauge action, while somewhat less effective at suppressing $\Delta_{\text{mix}}$, benefits from marginally weaker autocorrelations (see Appendix). Further investigation on the Borici-Creutz fermion and different choice of the staggered fermion (likes the stout smearing with different steps and smearing sizes) would be helpful to further suppress $\Delta_{\text{mix}}$ and worth studies in the future.
Furthermore, a preliminary study Lin and others (2025) of physical observable using clover valence fermions on our HISQ ensembles indicates that the discretization errors for quantities such as $f_{K}$ and $m_{\Omega}$ are suppressed compared to those from unitary clover fermion calculations Hu and others (2024); Hu et al. (2024). This suggests that the mixed-action setup can offer superior control over systematic uncertainty from continuum extrapolation, compared to the one with unitary valence-sea fermion action. Consequently, further investigations into a wider range of observables using this setup are warranted to identify optimal, cost-efficient strategies for lattice QCD calculations that maintain rigorous control over systematic errors.
@@ Acknowledgement
We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions Hu and others (2024), which are generated on HPC Cluster of ITP-CAS, IHEP-CAS and CSNS-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. The calculations were performed using the PyQUDA package Jiang et al. (2024) with QUDA Clark et al. (2010); Babich et al. (2011); Clark et al. (2016) through HIP programming model Bi et al. (2020). The numerical calculation were carried out on the ORISE Supercomputer, HPC Cluster of ITP-CAS and Advanced Computing East China Sub-center. This work is supported in part by National Key R&D Program of China No.2024YFE0109800, NSFC grants No. 12435002, 12525504, 12293060, 12293062, and 12447101, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. YSBR-101 and XDB34030303, and the science and education integration young faculty project of University of Chinese Academy of Sciences.
@@ IV.1 Autocorrelation Analysis
To assess the statistical independence of gauge configurations generated by the Hybrid Monte Carlo algorithm, we perform an autocorrelation analysis using the binning method following the approach in Ref. Du and others (2025a). The analysis focuses on four key observables: the average plaquette $\langle\text{Tr }U_{p}\rangle/3$, the topological charge $Q=\int d^{4}xq(x)$ with $q(x)=\frac{g^{2}}{32\pi^{2}}\epsilon_{\mu\nu\rho\sigma},\text{Tr},[F^{\mu\nu}(x)F^{\rho\sigma}(x)]$, the pion mass $m_{\pi}$, and the $\eta_{c}$ mass $m_{\eta_{c}}$.
For a given observable $O$, the normalized variance ratio $\sigma_{n}/\sigma_{1}$ is computed as a function of bin size $n$. First, the overall mean is calculated as $\bar{O}=\frac{1}{N}\sum_{i=1}^{N}O_{i}$, where $O_{i}$ denotes an individual measurement and $N$ is the total number of configurations. For a bin size $n$, the data is partitioned into $M=\left\lfloor\frac{N}{n}\right\rfloor$ bins, and the mean for the $k$-th bin is given by $\bar{O}_{n,k}=\frac{1}{n}\sum_{i=(k-1)n+1}^{kn}O_{i}$. The variance of these binned means is then computed as $\sigma_{n}^{2}(O)=\frac{1}{M}\sum_{k=1}^{M}\sigma_{n,k}^{2}(O)$ with $\sigma_{n,k}^{2}(O)=\frac{1}{M-1}(\bar{O}_{n,k}-\bar{O})^{2}$. To estimate the uncertainty of this variance, we calculate its standard error, $\mathrm{SE}(\sigma_{n}^{2})=\sqrt{\frac{1}{M(M-1)}\sum_{k=1}^{M}\left(\sigma_{n,k}^{2}-\sigma_{n}^{2}\right)^{2}}$.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@
The upper three panels of Fig. 5 illustrate the ratio $\sigma_{n}/\sigma_{1}$ as a function of bin size $n$ for all four observables, using data from the ensemble with $m_{\pi}\sim 0.3$ GeV at $a=0.108$ fm. These panels correspond to different molecular dynamics time steps: $\tau=0.5$ (upper left), $1.0$ (central left), and $2.0$ (lower left). As $\tau$ increases, the dependence of the statistical error on bin size weakens for all quantities.
The three panels in the middle present the cases for ensembles at finer lattice spacings ($a=0.09$, $0.07$, and $0.05$ fm from top to bottom), all with $\tau=2$. These plots suggest that the sensitivity to binning does not become significantly stronger at smaller lattice spacings.
The lower three panels show the cases with different gauge actions with increasing $|C_{P}|$, tree level Symanzik, tadpole improved Symanzik, and Iwasaki. We can see that the autocorrelation effect is somehow stronger for the pion mass with the Iwasaki gauge action.